# Improving $R(3,k)$ in just two bites

### v2

@@ 3 Preliminaries
Let $C=3\sqrt{20}$ and $\mathcal{D}$ be the event that the following hold for all distinct $v,w\in(V_{R}\cup V_{B})$.
$\big|F(v)-\log^{2}(n)\big|$ $\leq\varepsilon_{2}\log^{2}(n)$ (1)
$\big||N(v)|-pN\big|$ $\leq\varepsilon_{2}pN$ (2)
@@
$\big||N_{3}(v)|-pn\big|$ $\leq\varepsilon_{2}pn$ (4)
$|N_{3}(v)\cap N_{3}(w)|$ $\leq C\log^{3}(n)$ (5)
$|\pi_{B}(N_{3}(v))\cap\pi_{B}(N_{3}(w))|$ $\leq C\log^{3}(n)\mbox{ ~if $v,w\in V_{R}$}$ (6)
$|\pi_{R}(N_{3}(v))\cap\pi_{R}(N_{3}(w))|$ $\leq C\log^{3}(n)\mbox{ ~if $v,w\in V_{B}$}$ (7)
@@
Recall here that $1/n\ll\varepsilon_{2}\ll\varepsilon$ and therefore the conclusion that $\mathbb{P}(\mathcal{D})=1-o(1)$ is statement on asymptotic behavior as $n$ grows and $\varepsilon_{2}$ remains fixed.
Proof of Lemma 3.1.
[…]
Although the conclusion of the lemma takes all of the constants in the error to be the same (either $\varepsilon_{2}$ or $C$), due to dependencies between the involved quantities it is easier to prove these statements with slightly different constants. Thus we prove (1) and (2) with $\varepsilon_{2}^{\prime}=\varepsilon_{2}/3$, and we prove (3) with $C_{3}=\sqrt{20}$. Of course since $\varepsilon_{2}^{\prime}<\varepsilon_{2}$ and $C_{3}<C$, the claimed bounds will still hold.
[…]
$\mathbb{P}(\mathcal{D^{\prime}}^{c})\leq\mathbb{P}(A_{1}^{c})+\mathbb{P}(A_{2}^{c})+\mathbb{P}(A_{3}^{c})+\mathbb{P}(A_{4}^{c}\ |\ A_{1}\cap A_{2})+\mathbb{P}(A_{5}^{c}\ |\ A_{1}\cap A_{3})+\mathbb{P}(A_{6}^{c}\ |\ A_{1}\cap A_{2}\cap A_{3})+\mathbb{P}(A_{7}^{c}\ |\ A_{1}\cap A_{2}\cap A_{3})$
and analyze each of the terms on the right individually, starting with $\mathbb{P}(A_{1}^{c})$.
Since $G_{4}$ includes each vertex of $G_{1}$ independently, it is straightforward to check the expectation $\mathbb{E}(F(v))=\frac{n}{N^{2}}\cdot{}N=\log^{2}(n)$. Applying the Chernoff bound Theorem A.2 (here and for the remaining applications we are interested in random variables which are the sum of independent Bernoulli variables) with $t=c_{1}\log^{3/2}(n)$ shows that, for a fixed $v\in V_{R}\cup V_{B}$,
$\mathbb{P}(|F(v)-\log^{2}(n)|\geq c_{1}\log^{3/2}(n))\leq 2\exp\left(-\frac{c_{1}^{2}\log^{3}(n)}{4(\log^{2}(n))}\right)=O\left(\frac{1}{N^{4}}\right)$
where the last step holds for some $c_{1}$ a large constant. A union bound over $v$ shows that $\mathbb{P}(A_{1}^{c})=O(\frac{1}{N^{3}})$.
A similar strategy is effective for $A_{2}$. Again for a fixed $v$ we can calculate $\mathbb{E}(N(v))=\beta\sqrt{\frac{n}{\log^{3}(n)}}$, and applying Theorem A.2 now with $t=c_{2}\sqrt{\frac{n}{\log^{4}(n)}}$ shows
$\mathbb{P}\left(\left|N(v)-\beta\sqrt{\frac{n}{\log^{3}(n)}}\right|\geq c_{2}\sqrt{\frac{n}{\log^{4}(n)}}\right)\leq 2\exp\left(-\frac{c_{2}^{2}\sqrt{n}}{4(\log^{5/2}(n))}\right)=O\left(\frac{1}{N^{4}}\right)$
for some $c_{2}$ a large constant, and a union bound over $v$ shows $\mathbb{P}(A_{2}^{c})=O(\frac{1}{N^{3}})$.
For $A_{3}$ the expectation is much smaller since we assume $u\neq w$; now $\mathbb{E}(|N(v)\cap N(w)|)=Np^{2}=o(1)$. Then applying Theorem A.2 with $t=C_{3}\log n$ we see that for a fixed pair $v,w$
$\mathbb{P}\left(|N(v)\cap N(w)|\geq C_{3}\log n\right)\leq 2\exp\left(-\frac{C_{3}^{2}\log n}{4}\right)=O\left(\frac{1}{N^{5}}\right)$
for our choice of $C_{3}=\sqrt{20}$. A union bound over $v$ and $w$ shows that $\mathbb{P}(A_{3}^{c})=O(\frac{1}{N^{3}})$.
For $A_{4}$, since $(1+\varepsilon_{2}^{\prime})^{2}=(1+\varepsilon_{2}/3)^{2}\leq 1+\varepsilon_{2}$, it follows that $A_{1}\cap A_{2}\subseteq A_{4}$ and therefore $\mathbb{P}(A_{4}^{c}\ |\ A_{1}\cap A_{2})=0$. This deduction is one of the reasons why we prove (1) and (2) with modified constants.
The event $A_{5}$ considers two types of pairs $v,w$; ones where they both belong to  $V_{R}$ (or both to $V_{B}$) and pairs where, say, $v\in V_{R}$ and $w\in V_{B}$. In the first case, since $(1+\varepsilon_{2}^{\prime})(C_{3})\leq C$, (5) holds on $A_{1}\cap A_{3}$. In the second case $N_{3}(v)\cap N_{3}(w)$ is a  $(\frac{n}{N^{2}})$-random subset of the box $N(v)\times N(w)$ which has size $(1+o(1))p^{2}N^{2}$ on $A_{2}$. Computing the conditional expectation $\mathbb{E}(|N_{3}(v)\cap N_{3}(w)|\ |\ A_{1}\cap A_{3})=(1+o(1))\beta^{2}\log n$ and applying Theorem A.2 with $t=c_{5}\log n$ shows that
$\mathbb{P}\left(|N_{3}(v)\cap N_{3}(w)|\geq(2\beta^{2}+c_{5})\log n\right)\leq 2\exp\left(-\frac{c_{5}^{2}\log n}{4}\right)=O\left(\frac{1}{N^{5}}\right)$
for some sufficiently large fixed $c_{5}$ – again, $c_{5}=\sqrt{20}$ suffices. Observe then that $2\beta^{2}+\sqrt{20}\leq 3\sqrt{20}=C$. A union bound over the pairs $v\in V_{R},w\in V_{B}$ suffices to conclude that $\mathbb{P}(A_{5}^{c}\ |A_{1}\cap A_{3})=O(\frac{1}{N^{3}})$.
Finally we turn to $\mathbb{P}(A_{6}^{c}\ |\ A_{1}\cap A_{2}\cap A_{3})$ (the analysis for $A_{7}$ is entirely analogous). For $v,w\in V_{R}$, we decompose
$\pi_{B}(N_{3}(v))\cap\pi_{B}(N_{3}(w))\subseteq\left(\pi_{B}(N_{3}(v)\cap N_{3}(w))\cup\bigg(\pi_{B}(N_{3}(v)\setminus N_{3}(w))\cap\pi_{B}(N_{3}(w)\setminus N_{3}(v))\bigg)\right).$ (8)
If $A_{1}\cap A_{3}$ occurs, then deterministically
$|\pi_{B}(N_{3}(v)\cap N_{3}(w))|\leq(1+\varepsilon_{2}^{\prime})C_{3}\log^{3}(n)<2\sqrt{20}\log^{3}(n).$
For the second term, let $Z_{v,w}=|\pi_{B}(N_{3}(v)\setminus N_{3}(w))\cap\pi_{B}(N_{3}(w)\setminus N_{3}(v))|$. To compute $\mathbb{E}(Z_{v,w})$, consider a fixed element $b\in V_{B}$, which lies in the above set if there are $x\in N(v)\setminus N(w)$ and $y\in N(w)\setminus N(v)$ with $(x,b)\in V(G_{3})$ and $(y,b)\in V(G_{3})$. Using now that $N(w)\setminus N(v)$ and $N(v)\setminus N(w)$ are disjoint, this occurs with probability $(1-(1-n/N^{2})^{|N(v)\setminus N(w)|})(1-(1-n/N^{2})^{|N(w)\setminus N(v)|})$. Having conditioned on $A_{2}$ and $A_{3}$, the intersection $N(v)\cap N(w)$ is small and therefore it still holds that $|N(v)\setminus N(w)|=(1+o(1))pN$, and likewise for $|N(w)\setminus N(v)|$. Using the large $n$ approximation for $(1-n/N^{2})^{(1+o(1))pN}$$(1+o(1))(1-\frac{pn}{N})$, we see
$\mathbb{E}(Z_{v,w}\ |\ A_{1}\cap A_{2}\cap A_{3})=(1+o(1))\beta^{2}N\frac{\log^{5}(n)}{n}\leq\log^{3}(n),$
for $n$ sufficiently large. Furthermore, this intersection is binomially distributed, as for different $b$ these events are independent.
Now, applying Theorem A.2 with $t=c_{6}\log^{2}(n)$ implies
$\mathbb{P}\left(Z_{v,w}\geq\log^{3}(n)+c_{6}\log^{2}(n)\right)\leq 2\exp\left(-\frac{c_{6}^{2}\log n}{4}\right)=O\left(\frac{1}{N^{5}}\right)$
for some sufficiently large $c_{6}>0.$ Combining the bounds obtained on $Z_{v,w}$ and $|\pi_{B}(N_{3}(v)\cap N_{3}(w))|$ we see that
$|\pi_{B}(N_{3}(v))\cap\pi_{B}(N_{3}(w))|\leq|\pi_{B}(N_{3}(v)\cap(N_{3}(w))|+Z_{v,w}\leq 3\sqrt{20}\log^{3}(n)=C\log^{3}(n)$
with high probability – and as before, $\mathbb{P}(A_{6}^{c}\ |\ A_{1}\cap A_{2}\cap A_{3})=O(\frac{1}{N^{3}})$ by the union bound. Likewise, by symmetry $\mathbb{P}(A_{7}^{c}\ |\ A_{1}\cap A_{2}\cap A_{3})=O(\frac{1}{N^{3}}).$
Summing the bounds obtained above for each $A_{i}$ shows that $\mathbb{P}(\mathcal{D^{\prime}}^{c})=O(\frac{1}{N^{3}})$.
We are now ready to transfer this analysis back to our model where $V(G_{2})$ is selected randomly from all $n$-vertex subsets. As remarked at the start of the proof,  $\mathbb{P}(\mathcal{D^{\prime}}^{c}\ |\ v(G^{\prime})=n)=\mathbb{P}(\mathcal{D}^{c})$. One can also observe that $\mathbb{P}(|V(G_{3})|=n)=\Omega(\frac{1}{N^{2}})$, since $|V(G_{3})|$ as a random variable takes on $N^{2}+1$ discrete values and $n$ is the most probable. Putting these together,
$\mathbb{P}(\mathcal{D}^{c})\leq\frac{\mathbb{P}(\mathcal{D}^{\prime c})}{\mathbb{P}(|V(G_{3})|=n)}=O\left(\frac{1}{N}\right)$
and we are done. ∎
Given $I\subseteq V(G)$ with $|I|=k$ and a vertex $v\in V_{R}\cup V_{B}$, we let
$X_{v}(I)=\varphi(I)\cap X_{v},\hskip 28.45274ptX_{v}^{+}(I)=\varphi(I)\cap X_{v}^{+}.$
@@
‘closed’ and ‘closed+’ pairs, while the pairs in $O(I)={\varphi(I)\choose 2}\setminus C(I)$ and $O^{+}(I)={\varphi(I)\choose 2}\setminus C^{+}(I)$ are ‘open’ and ‘open+’, respectively. By definition, we have $C^{+}(I)\subseteq C(I)$ and $O(I)\subseteq O^{+}(I)$. As $G_{2}$ is triangle-free, all edges in $G_{2}$ are in $O^{+}(I)$. The main challenge then is to understand the set $C^{+}(I)$ for various $I$, which we do through a series of lemmas. Due to the greater symmetry it is easier to consider $C(I)$, though, and all lemmas in this section contend with $C(I)$.
@@
and partition $V_{R}\cup V_{B}$ by size as
$H_{I}$ $=\left\{v\in V_{R}\cup V_{B}:|X_{v}(I)|>t_{1}\right\},$
@@
$M_{I}$ $=\left\{v\in V_{R}\cup V_{B}:t_{3}<|X_{v}(I)|<t_{2}\right\},$
$S_{I}$ $=\left\{v\in V_{R}\cup V_{B}:|X_{v}(I)|<t_{3}\right\}.$
@@
$k=|I|\geq\sum_{v\in L}|X_{v}(I)|-\sum_{\{v,w\}\in{L\choose 2}}|X_{v}(I)\cap X_{w}(I)|\geq|L|n^{1/4+\varepsilon}-\binom{|L|}{2}3\sqrt{20}\log^{3}(n).$
@@
$\sum_{v\in L_{I}}\binom{|X_{v}(I)|}{2}\leq\sum_{v\in L_{I}}\frac{|X_{v}(I)|t_{1}}{2}=o(k^{2})$
since $t_{1}=o(k)$, and the result follows for $n$ sufficiently large. ∎
Let us now look at the sets $X_{v}(I)$ of the next smaller size – those with $v\in M_{I}$. The following is a weaker version of a lemma in [campos2025], showing that the contribution of closed pairs from $M_{I}$ is $o(k^{2})$, with high probability.
@@
Suppose that $Y_{I}>kn^{1/4-\varepsilon}.$ Let $B\subseteq M_{I}$ be minimal so that $Y=\sum_{v\in B}|X_{v}(I)|>kn^{1/4-\varepsilon}$. Note that $Y<2kn^{1/4-\varepsilon}$ (since each $|X_{v}(I)|\leq n^{1/4+\varepsilon}$) and that $|B|\leq n^{-2\varepsilon}Y\leq 2kn^{1/4-3\varepsilon}.$ Let $B_{R}=B\cap V_{R}$ and $B_{B}=B\cap V_{B}$. Then $Y$ is the number of edges in the union of two bipartite graphs $\Gamma_{R}$ and $\Gamma_{B}$ on $V(\Gamma_{R})=(B_{R},\varphi(I))$, $V(\Gamma_{B})=(B_{B},\varphi(I))$ where
@@
Now, consider a fixed $\tilde{B}=\tilde{B}_{R}\cup\tilde{B}_{B}\subseteq(V_{R}\cup V_{B})$ of size $2kn^{1/4-3\varepsilon}$ and $I\subseteq V(G)$ of size $k$. Suppose $Z\sim Bin(|\tilde{B}|k,p)$. The probability that there are at least $\frac{kn^{1/4-\varepsilon}}{3\log^{2}(n)}$ edges in the union of the two bipartite subgraphs on $(\tilde{B}_{R},\pi_{R}(I))$ in $G_{R}$ and on $(\tilde{B}_{B},\pi_{B}(I))$ in $G_{B}$ is at most (for sufficiently large $n$)
@@
$\binom{n}{k}\binom{2N}{2kn^{1/4-3\varepsilon}}\leq\exp\big(k\log n+2kn^{1/4-3\varepsilon}\log n\big)=\exp\big(o(n^{3/4-2\varepsilon})\big),$
and that $B\subseteq\tilde{B}$ for some $\tilde{B}$, a union bound over all possible choices of $\tilde{B}$ and $I$ shows that with high probability none of the potential choices of $(B,\varphi(I))$ yield graphs with as many edges as required above.
Letting $\mathcal{O}$ denote the existence of such an overfull $(B,\varphi(I))$, we conclude that the probability that the high probability conclusion of the lemma is violated is at most $\mathbb{P}(\mathcal{O}^{c}\cup\mathcal{D}^{c})=o(1).$
@@
For a pair $e\in{F(w)\cap\varphi(I)\choose 2}$ for some $w\in V_{R}\cup V_{B}$, $e\in{X_{v}(I)\choose 2}$ for any $v$ with $vw\in E(G_{R})\cup E(G_{B})$. Assuming $\mathcal{D}$, we have $|F(w)\cap\varphi(I)|<2\log^{2}(n)$ for all $w$. As the sets $F(w)\choose 2$} are disjoint, and every vertex in $\varphi(I)$ is in exactly two fibers, there are fewer than $\frac{2k}{2\log^{2}{n}}2\log^{4}(n)=2k\log^{2}(n)$ such pairs.
For a fixed pair $e\in\binom{\varphi(I)}{2}\setminus\bigcup_{w}{F(w)\choose 2}$, the probability that $e$ lies in the union $\bigcup_{v}\binom{X_{v}(I)}{2}$ (regardless of whether or not $v\in S_{I}$) is at most
$2\bigl(1-(1-p^{2})^{N}\bigr)\;\leq\;2p^{2}N\;=\;\frac{2\beta^{2}}{\log n}.$
@@
indexed by $v\in(V_{R}\setminus\pi_{R}(I))\cup(V_{B}\setminus\pi_{B}(I))$}. For fixed $\varphi$, these variables depend only on $N(v)$, which is determined by the choices of $G_{R}$ and $G_{B}$. They are mutually independent because we are only considering those where $v\not\in\pi_{R}(I)\cup\pi_{B}(I)$. As the union over such $v$, $Z$ is a deterministic function of the $Y_{v}$ and changing the value of a single such $Y_{v}$ only alters the set $\binom{X_{v}(I)}{2}$ therefore changing the value of $Z$ by at most
@@
As there are $|S_{I}\setminus\pi_{R}(I)\cup\pi_{B}(I)|\leq 2N$ random variables in our family, McDiarmid’s Inequality (Theorem A.4) yields, for any $t>0$,
$\mathbb{P}\left(Z\geq\mathbb{E}(Z)+t\right)\leq\exp\left(-\,\frac{t^{2}}{N\,n^{8\varepsilon}}\right).$
@@
$\leq\exp\left(-\,\frac{\bigl(\frac{1}{\sqrt{\log n}}\binom{k}{2}\bigr)^{2}}{N\,n^{8\varepsilon}}\right)$
@@
for all sufficiently large $n$. Combining this with (10), we find that, for a fixed $I$, the whole quantity
@@
Finally, we turn to the largest of the $X_{v}(I)$ – those with $v\in H_{I}$. Here is where the biggest potential contributions to the closed pairs come from.
@@
If $\mathcal{D}$ of Lemma 3.1 occurs, and $n$ is sufficiently large, then all $I\subseteq V(G)$ with $|I|=k$ satisfy
@@
From this and the fact (from $\mathcal{D}$) that all vertices in $V_{R}$ and $V_{B}$ have degree at most $(1+\varepsilon_{2})pN\leq(\beta+\varepsilon_{2})\sqrt{\frac{n}{\log^{3}(n)}}$ we have
$\sum_{v\in H_{I}\cap V_{R}}|\pi_{R}(X_{v}(I))|\leq|H_{I}|(1+\varepsilon_{2})pN=o(k)$ and $\sum_{v\in H_{I}\cap V_{B}}|\pi_{B}(X_{v}(I))|=o(k).$ (17)
@@
Order the vertices of $H_{I}\cap V_{R}$ as $v_{1},\dots,v_{t}$, and set $x_{v_{i}}=|\pi_{B}(X_{v_{i}}(I))|$. We may assume that $x_{v_{1}}$ is the maximum of the $x_{v_{i}}$. Note that $x_{v_{1}}\leq|N_{3}(v_{1})|\leq(1+\varepsilon_{2})pn$. By the convexity of the binomial coefficients, we have for sufficiently large $n$
@@
$|\pi_{R}(I)|$ $\leq\left|\pi_{R}\left(\bigcup X_{v_{i}}(I)\right)\right|+\left|\pi_{R}\left(\varphi(I)\setminus\left(\bigcup X_{v_{i}}(I)\right)\right)\right|$
$\leq\left|\pi_{R}\left(\bigcup X_{v_{i}}(I)\right)\right|+\left|\varphi(I)\setminus\left(\bigcup X_{v_{i}}(I)\right)\right|$
@@
We now turn to the crucial part of Theorem 1.3, showing that vertex sets of size $k$ are not independent. This is first and only time that it is important to remove edges from monochromatic triangles by considering $C^{+}(I)$. Up until this time, we could have omitted all edges in monochromatic triangles by using $C(I)$ instead.
@@
    First, let $F_{0}=(V_{R}\cup V_{B})\setminus(\pi_{R}(I)\cup\pi_{B}(I))$. For all vertices $v\in F_{0}$, reveal all neighbors and non-neighbors in $\pi_{R}(I)$ and $\pi_{B}(I)$ in $G_{R}$ and $G_{B}$ respectively. These, in turn, reveal edges within $G_{2}$.
@@
    Let $F_{1}=\{v\in\pi_{R}(I)\cup\pi_{B}(I):|F(v)\cap I|>\log n\}$. For all vertices in $F_{1}$, reveal all neighbors and non-neighbors in $\pi_{R}(I)$ and $\pi_{B}(I)$ in $G_{R}$ and $G_{B}$ respectively. These, in turn, reveal edges within $G_{2}$.
@@
    $|X_{v}(I)|<(1+\varepsilon_{2})(\log^{2}n)\frac{t_{2}}{\log n}+(\log n)(1+\varepsilon_{2})pN\ll t_{1},$
@@
    For all vertices in $F_{2}$, reveal all neighbors and non-neighbors in $\pi_{R}(I)$ and $\pi_{B}(I)$ in $G_{R}$ and $G_{B}$ respectively. These, in turn, reveal edges within $G_{2}$.

### v3

@@ 3 Controlling the edge deletions
Let $C=100$ and $\mathcal{D}$ be the event that the following hold for all distinct $v,w\in(V_{R}\cup V_{B})$.
$\big||F(v)|-\log^{2}(n)\big|$ $\leq\varepsilon_{2}\log^{2}(n)$ (1)
$\big||N(v)|-p{m}\big|$ $\leq\varepsilon_{2}p{m}$ (2)
@@
$\big||N_{v}|-pn\big|$ $\leq\varepsilon_{2}pn$ (4)
$|N_{v}\cap N_{w}|$ $\leq C\log^{3}(n)$ (5)
$|\pi_{B}(N_{v})\cap\pi_{B}(N_{w})|$ $\leq 2C\log^{3}(n)\mbox{ ~if $v,w\in V_{R}$}$ (6)
$|\pi_{R}(N_{v})\cap\pi_{R}(N_{w})|$ $\leq 2C\log^{3}(n)\mbox{ ~if $v,w\in V_{B}$}$ (7)
@@
Proof sketch.
We only sketch the proof here, the details are left to the reader. We show that each statement fails with probability $o(1)$. Statements (2) and (3) are standard results about degrees and co-degrees in the random graphs $G_{R}$ and $G_{B}$. Statements (1), (4) and (5) follow from Chernoff bounds (Theorem A.2) for $\pi(V(G))$ intersecting the various subsets of $V_{R}\times V_{B}$ determined by $G_{R}$ and $G_{B}$, taking the union over all choices of $v$ and $w$. Note that the choice of $\pi$ being injective only strengthens the concentration around the various means. Finally, for (6) and (7), we first use (5) for a possible intersection, and then we use Chernoff bounds for the hypergeometric distribution for intersections of two random subsets of size $(1+\varepsilon_{2})pn$ of a set of size m{m}. ∎
One main difficulty in the proof of Theorem 1.3 is to analyze the edges in $\tilde{E}(G)\setminus E(G)$ removed to make the graph triangle-free. Every such edge is between the two end-vertices of a monochromatic path of length $2$, so we study such vertex pairs in $N_{v}$ for $v\in V_{R}\cup V_{B}$. Given $I\subseteq V(G)$ with $|I|=k$ and a vertex $v\in V_{R}\cup V_{B}$, we let
$X_{v}(I)=I\cap N_{v},\hskip 28.45274ptX_{v}^{+}(I)=I\cap N_{v}^{+}.$
@@
‘closed’ and ‘closed+’ pairs111this nomenclature is inherited from the triangle-free process, where pairs are closed when we can not add an edge without creating a triangle, while the pairs in $O(I)={\pi(I)\choose 2}\setminus C(I)$ and $O^{+}(I)={\pi(I)\choose 2}\setminus C^{+}(I)$ are ‘open’ and ‘open+’, respectively. Edges in $\tilde{E}(G)\setminus E(G)$ deleted due to monochromatic triangles are in $C^{+}(I)\subseteq C(I)$, whereas edges deleted due to $2$-colored triangles may be anywhere in $C(I)$.
Thus, we want to understand $C(I)$ and $C^{+}(I)$ for various $I$, which we do through a series of lemmas. Due to the greater symmetry it is easier to consider only the larger sets $C(I)$, though, and all lemmas in this section contend with $C(I)$.
@@
and partition $V_{R}\cup V_{B}$ by size (huge, large, medium, small) as
$H_{I}$ $=\left\{v\in V_{R}\cup V_{B}:t_{1}<|X_{v}(I)|\leq k~\right\},$
@@
$M_{I}$ $=\left\{v\in V_{R}\cup V_{B}:t_{3}<|X_{v}(I)|\leq t_{2}\right\},$
$S_{I}$ $=\left\{v\in V_{R}\cup V_{B}:0~\leq|X_{v}(I)|\leq t_{3}\right\}.$
@@
$k=|I|\geq\sum_{v\in L}|X_{v}(I)|-\sum_{\{v,w\}\in{L\choose 2}}|X_{v}(I)\cap X_{w}(I)|\geq|L|n^{1/4+\varepsilon}-\binom{|L|}{2}C\log^{3}(n).$
@@
$\sum_{v\in L_{I}}\binom{|X_{v}(I)|}{2}\leq\sum_{v\in L_{I}}\frac{|X_{v}(I)|t_{1}}{2}=\frac{t_{1}}{2}\sum_{v\in L_{I}}|X_{v}(I)|=o(k^{2}),$
and the result follows for $n>n_{0}$. ∎
Let us now look at the sets $X_{v}(I)$ of the next smaller size – those with $v\in M_{I}$. The following is a weaker version of a lemma in [8], showing that the contribution of closed pairs from $M_{I}$ is $o(k^{2})$, with high probability.
@@
Suppose that $Y_{I}>kn^{1/4-\varepsilon}.$ Let $B\subseteq M_{I}$ be minimal so that $Y=\sum_{v\in B}|X_{v}(I)|>kn^{1/4-\varepsilon}$. Note that $Y<2kn^{1/4-\varepsilon}$ (since each $|X_{v}(I)|\leq n^{1/4+\varepsilon}$) and that $|B|\leq n^{-2\varepsilon}Y\leq 2kn^{1/4-3\varepsilon}.$ Let $B_{R}=B\cap V_{R}$ and $B_{B}=B\cap V_{B}$. Then $Y$ is the number of edges in the union of two bipartite graphs $\Gamma_{R}$ and $\Gamma_{B}$ on $V(\Gamma_{R})=(B_{R},\pi(I))$, $V(\Gamma_{B})=(B_{B},\pi(I))$ where
@@
Now, consider a fixed $\tilde{B}=\tilde{B}_{R}\cup\tilde{B}_{B}\subseteq(V_{R}\cup V_{B})$ of size $2kn^{1/4-3\varepsilon}$ and $I\subseteq V(G)$ of size $k$. Suppose $Z\sim Bin(|\tilde{B}|k,p)$. The probability that there are at least $\frac{kn^{1/4-\varepsilon}}{3\log^{2}(n)}$ edges in the union of the two bipartite subgraphs on $(\tilde{B}_{R},\pi_{R}(I))$ in $G_{R}$ and on $(\tilde{B}_{B},\pi_{B}(I))$ in $G_{B}$ is at most (for $n>n_{0}$)
@@
$\binom{n}{k}\binom{2{m}}{2kn^{1/4-3\varepsilon}}\leq\exp\big(k\log n+2kn^{1/4-3\varepsilon}\log n\big)=\exp\big(o(n^{3/4-2\varepsilon})\big),$
and that $B\subseteq\tilde{B}$ for some $\tilde{B}$, a union bound over all possible choices of $\tilde{B}$ and $I$ shows that with high probability none of the potential choices of $(B,\pi(I))$ yield graphs with as many edges as required above.
Letting $\mathcal{O}$ denote the existence of such an overfull $(B,\pi(I))$, we conclude that the probability that the high probability conclusion of the lemma is violated is at most $\mathbb{P}(\mathcal{O}^{c}\cup\mathcal{D}^{c})=o(1).$
@@
For a pair $e\in{F(w)\cap\pi(I)\choose 2}$ for some $w\in V_{R}\cup V_{B}$, $e\in{X_{v}(I)\choose 2}$ for any $v$ with $vw\in E(G_{R})\cup E(G_{B})$. Assuming $\mathcal{D}$, we have $|F(w)\cap\pi(I)|<2\log^{2}(n)$ for all $w$. As the sets $F(w)\choose 2$} are disjoint, and every vertex in $\pi(I)$ is in exactly two fibers, there are fewer than $\frac{2k}{2\log^{2}{n}}2\log^{4}(n)=2k\log^{2}(n)$ such pairs.
For a fixed pair $e\in\binom{\pi(I)}{2}\setminus\bigcup_{w}{F(w)\choose 2}$, the probability that $e$ lies in the union $\bigcup_{v}\binom{X_{v}(I)}{2}$ (regardless of whether or not $v\in S_{I}$) is at most
$2\bigl(1-(1-p^{2})^{m}\bigr)\;\leq\;2p^{2}{m}\;=\;\frac{2\beta^{2}}{\log n}.$
@@
indexed by $v\in(V_{R}\setminus\pi_{R}(I))\cup(V_{B}\setminus\pi_{B}(I))$}. For fixed $\pi$, these variables depend only on $N(v)$, which is determined by the choices of $G_{R}$ and $G_{B}$. They are mutually independent because we are only considering those where $v\not\in\pi_{R}(I)\cup\pi_{B}(I)$. As the union over such $v$, $Z$ is a deterministic function of the $Y_{v}$ and changing the value of a single such $Y_{v}$ only alters the set $\binom{X_{v}(I)}{2}$ therefore changing the value of $Z$ by at most
@@
As there are $|S_{I}\setminus\pi_{R}(I)\cup\pi_{B}(I)|\leq 2{m}$ random variables in our family, McDiarmid’s Inequality (Theorem A.4) yields, for any $t>0$,
$\mathbb{P}\left(Z\geq\mathbb{E}(Z)+t\right)\leq\exp\left(-\,\frac{t^{2}}{{m}\,n^{8\varepsilon}}\right).$
@@
$\leq\exp\left(-\,\frac{\bigl(\frac{1}{\sqrt{\log n}}\binom{k}{2}\bigr)^{2}}{{m}\,n^{8\varepsilon}}\right)$
@@
for $n>n_{0}$. Combining this with (9), we find that, for a fixed $I$, the whole quantity
@@
Finally, we turn to the largest of the $X_{v}(I)$ – those with $v\in H_{I}$. Here is where the only significant contributions to the closed pairs arise. Ultimately, the independence of $I$ is determined by the behavior of $G_{R}$ and $G_{B}$ in $\pi_{R}(I)$ and $\pi_{B}(I)$, so we focus on these projected sets here.
@@
If $\mathcal{D}$ of Lemma 3.1 occurs, and $n>n_{0}$, then all $I\subseteq V(G)$ with $|I|=k$ satisfy
@@
From this and the fact (from $\mathcal{D}$) that all vertices in $V_{R}$ and $V_{B}$ have degree at most $(1+\varepsilon_{2})p{m}\leq(\beta+\varepsilon_{2})\sqrt{\frac{n}{\log^{3}(n)}}$ we have
$\sum_{v\in H_{I}\cap V_{R}}|\pi_{R}(X_{v}(I))|\leq|H_{I}|(1+\varepsilon_{2})p{m}=o(k)$ and $\sum_{v\in H_{I}\cap V_{B}}|\pi_{B}(X_{v}(I))|=o(k).$ (16)
@@
Order the vertices of $H_{I}\cap V_{R}$ as $v_{1},\dots,v_{t}$, and set $x_{v_{i}}=|\pi_{B}(X_{v_{i}}(I))|$. We may assume that $x_{v_{1}}$ is the maximum of the $x_{v_{i}}$. Note that $x_{v_{1}}\leq|N_{3}(v_{1})|\leq(1+\varepsilon_{2})pn$. By the convexity of the binomial coefficients, we have for $n>n_{0}$
@@
$|\pi_{R}(I)|$ $\leq\left|\pi_{R}\left(\bigcup X_{v_{i}}(I)\right)\right|+\left|\pi_{R}\left(\pi(I)\setminus\left(\bigcup X_{v_{i}}(I)\right)\right)\right|$
$\leq\left|\pi_{R}\left(\bigcup X_{v_{i}}(I)\right)\right|+\left|\pi(I)\setminus\left(\bigcup X_{v_{i}}(I)\right)\right|$
@@
We now turn to the crucial part of Theorem 1.3, showing that vertex sets of size $k$ are not independent. The only time that it is important to remove edges from monochromatic triangles by considering only $C^{+}(I)$ is in one place in the proof of Lemma 4.1 below. For all previous arguments, we could have omitted all edges in monochromatic triangles since we bound $C(I)$ instead.
@@
    First, let $F_{0}=(V_{R}\cup V_{B})\setminus(\pi_{R}(I)\cup\pi_{B}(I))$. For all vertices $v\in F_{0}$, reveal all neighbors and non-neighbors in $\pi_{R}(I)$ and $\pi_{B}(I)$ in $G_{R}$ and $G_{B}$ respectively. These, in turn, reveal edges in $\tilde{E}(G)$.
@@
    Let $F_{1}=\{v\in\pi_{R}(I)\cup\pi_{B}(I):|F(v)\cap I|>\log n\}$. For all vertices in $F_{1}$, reveal all neighbors and non-neighbors in $\pi_{R}(I)$ and $\pi_{B}(I)$ in $G_{R}$ and $G_{B}$ respectively. These, in turn, reveal edges in $\tilde{E}(G)$.
@@
    $|X_{v}(I)|<(1+\varepsilon_{2})(\log^{2}n)\frac{t_{2}}{\log n}+(\log n)(1+\varepsilon_{2})p{m}\ll t_{1},$
@@
    For all vertices in $F_{2}$, reveal all neighbors and non-neighbors in $\pi_{R}(I)$ and $\pi_{B}(I)$ in $G_{R}$ and $G_{B}$ respectively. These, in turn, reveal edges in $\tilde{E}(G)$.